twenty-three less than 4 times a number is twenty-one.What is the number?
Answer by rm29924(97)

let x= a number
twenty-three less than 4 times a number is twenty-one
twenty-three less means we will subtract 23 from 4 times a number
4 times a number - 23=21
4x-23=21
add 23 to both sides
4x=44
divide both sides by 4 to get x by itself
x=11
check answer
4x-23=21
4(11)-23=21
44-23=21
21=21